OpenStudy (anonymous):

equation for this situation:the cost,c, in dollars of a car rental is equal to 10dollars plus 1/2 the miles driven,m determine the cost of the car rental if the number of miles driven is 12.

OpenStudy (dumbcow):

c = 10 + m/2 m=12 c = 10 + 6 = 16